OpenStudy (anonymous):

there are 25 students in a classroom and everyone gets really excited and high fives everyone esle. How many high fives are exvahnged if every student gives a high five to every one of his/her classmates ?

OpenStudy (anonymous):

exchanged*

OpenStudy (binary3i):

\[\left(\begin{matrix}25 \\ 2\end{matrix}\right)\] factorial(25)/{factorial(2)*factorial(23)}=25*12=300

OpenStudy (anonymous):

where did you get 12 from ?